# Quantum Computing for Energy Correlators

### v1

@@ II.3.2 Hadamard test
After preparing the interacting ground state, we introduce an ancilla qubit and perform the Hadamard test: We first apply a Hadamard gate to an ancilla qubit, then apply a controlled-unitary gate for the operator insertions in the real-time correlator $\langle{\rm vac}|\mathcal{O}^{\dagger}T_{0i}(t_{1})T_{0j}(t_{2})\mathcal{O}|{\rm vac}\rangle$ [^5] We suppress the spatial dependence here for convenience. The time dependence is simulated as shown Eq. (II.2)., and finally apply another Hadamard gate and measure the ancilla qubit, as shown in Fig. 4a The probability of the ancilla qubit being in $|0\rangle$ state is given by
@@
The source ($\mathcal{O}$), sink ($\mathcal{O}^{\dagger}$) and the energy flux $T_{0i}$ operators are not unitary, making their direct implementation on the quantum circuit hard. Here we assume the source and sink operators are Hermitian (the energy flux operator $T_{0i}$ is Hermitian) and propose three different methods to implement them.
@@
    Taylor expansion of exponential. Suppose that $\mathcal{A}$ is a Hermitian operator and we have to obtain $\mathcal{A}\ket{\psi}$, we can implement two quantum circuits where one circuit implements $e^{i\theta\mathcal{A}}$ gate and the other $e^{-i\theta\mathcal{A}}$. The difference of the two circuits gives
@@
    In this way, we can implement $\mathcal{A}\ket{\psi}$ up to a small error when $\theta$ is small. In our case, we have to implement four different non-unitary operators, $\{\mathcal{O}^{\dagger}=\mathcal{O},\,T_{0i}(t_{1}),\,T_{0j}(t_{2}),\,\mathcal{O}\}$. In order to obtain the real-time correlator $\langle{\rm vac}|\mathcal{O}T_{0i}(t_{1})T_{0j}(t_{2})\mathcal{O}|{\rm vac}\rangle$, we should run 16 different quantum circuits, in each of which we implement the operator of the form $e^{is_{1}\theta\mathcal{O}}e^{is_{2}\theta T_{0i}(t_{1})}e^{is_{3}\theta T_{0j}(t_{2})}e^{is_{4}\theta\mathcal{O}}$ for one sign combination $s_{i}=\pm$. There are totally 16 sign combinations, corresponding to the 16 quantum circuits. By taking a proper linear combination of the results on the 16 quantum circuits, we find
@@
    where the $2\times 2$ block structure is associated with the additional ancilla qubit. After the implementation of $U_{\mathcal{A}}$, the ancilla qubit is measured. When the measurement returns |$\ket{0}$, the correct operator is applied, i.e., $\mathcal{A}$. The success probability of applying $\mathcal{A}$ to |$\ket{\psi}$ is given by $P_{s}=\|\frac{\mathcal{A}}{\|\mathcal{A}\|}|\psi\rangle\|$, which is equal to the expectation value $\langle{\psi}|\frac{A^{2}}{\|A\|^{2}}|{\psi}\rangle$.
@@
    Linear combination of unitaries (LCU). The $\mathcal{O}$ and $T_{0i}$ operators of our interest can be written as sums of Pauli strings that are unitary operators. Therefore, we can apply the method of linear combination of unitaries (LCU) [141]. To implement an operator $\mathcal{A}$ given by a linear combination of unitary operators $\mathcal{A}=\sum_{j=1}^{N}\kappa_{j}U_{j}$ with $\kappa_{j}$ real numbers, we can add $m$ additional ancilla qubits, where $m$ is given by $\lceil\log_{2}N\rceil$, and implement the quantum circuit shown in Fig. 5. The $V$ gate is given by a unitary that implements
@@
    Figure 5: LCU quantum circuit.
    Analogous to the diluted operator case, the correct implementation of $\mathcal{A}$ is obtained if all the $m$ ancilla qubits are measured to be |$\ket{0}$. The success probability is given by
    $P_{s}=\frac{\|\sum_{j=1}^{N}\kappa_{j}U_{j}|\psi\rangle\|}{\|\kappa\|_{1}}\,.$ (22)
@@ III.1 Hamiltonian on honeycomb lattice
with $a,b,c\in\{1,2,3\}$ for SU(2) adjoint indices. $\varepsilon^{abc}$ represents the structure constant of the non-Abelian group and is the Levi-Civita tensor for the SU(2) case..
@@
Electric fields are generators of time-independent gauge transformation. On the honeycomb lattice, since electric fields live on links, they can generate gauge transformations on either the red or the black vertices. So we distinguish these two types by the subscript $R$ for red and $B$ for black. The gauge transformation is given by
@@
Because of the Peter-Weyl theorem, we can work in the electric basis where the basis states are specified by three quantum numbers $|jm_{R}m_{B}\rangle$, analogous to the total angular momentum quantum number $j$ and two third components $m_{R}$ and $m_{B}$ for the red and black vertices, respectively. In the electric basis, the matrix elements of the electric fields are given by [143, 144, 145]
@@
where $|j\rangle$ and $|J\rangle$ denote the states on a link where the electric field squared acts. The matrix elements of the electric energy are diagonal in the electric basis and they are independent of whether the red or the black vertex is used. The matrix elements of the magnetic energy (the honeycomb plaquette) are given by [146, 147, 142] (see Refs. [148, 149, 150, 151, 152] for the square plaquette case)
@@ III.2 Energy flux operator
To construct energy flow operator, we need specific components of the stress-energy tensor. In particular, the energy flux operator is given as
@@
which we need to integrate in order to build the energy flow operator as in Eq. (1). Here $k,\ell$ denote spatial indices.
On a honeycomb lattice, the position ${\boldsymbol{x}}$ of the $T_{0k}(\boldsymbol{x})$ operator is given by the red vertex as shown in the Fig. 6, which can be labeled by $(i,j)$. The natural directions are no longer $\hat{x}$ and $\hat{y}$, but rather $\hat{e}_{1}$, $\hat{e}_{2}$ and $\hat{e}_{3}$. We can use the lattice discretized version of Eq. (36) and define three energy flux operators corresponding to the three electric field operators $E_{Ri}^{a}$ along the three directions $\hat{e}_{i}$ (see Appendix A). These energy fluxes point from red vertices to red vertices. Therefore, to indicate both the spatial position of the energy flux operator and the direction of the energy flow, we find it more convenient to introduce a notation to denote energy flux from the red vertex $(i,j)$ pointing towards $(i^{\prime},j^{\prime})$ as
@@
From the expression given in Eq. (36), we are able to express $T_{0,(i,j)\to(i^{\prime},j^{\prime})}$ for our $2+1D$ SU(2) on the honeycomb lattice in terms of the electric field and the plaquette operator. We carry out this exercise in the Appendix A. However, the energy flux operator obtained by discretizing Eq. (36) is a lattice bare operator and probably affected by the boundary condition. In general one needs to tame UV divergences and renormalize it in the continuum limit, since the lattice breaks Lorentz symmetry and the stress-energy tensor obtained in the continuum may not satisfy the conservation equation $\partial_{\mu}T^{0\mu}=0$ on the lattice. Since energy correlators describe how energy flows in the system, it is important to use a lattice energy flux operator that exactly conserves energy in the lattice calculation. Such a physical and IR-finite operator can be constructed by considering the Hamiltonian density at lattice position $(i,j)$ in the Heisenberg picture $H_{ij}(t)=e^{iHt}H_{ij}e^{-iHt}$. Taking time derivatives gives
$\frac{{\rm d}H_{ij}(t)}{{\rm d}t}=i[H,H_{ij}]\,.$ (38)
@@
On a honeycomb lattice, six possible energy flux directions $(i^{\prime},j^{\prime})$ for a given $(i,j)$ are $(i^{\prime},j^{\prime})\in\{(i,j+1),(i+1,j),(i+1,j-1),(i,j-1),(i-1,j),(i-1,j+1)\}$, as shown in Fig. 7, where the current $T_{0,(i,j)\to(i^{\prime},j^{\prime})}$ is nonvanishing. In Appendix A, we explicitly demonstrate that the lattice bare energy flux operator is identical to Eq. (40) up to boundary terms and UV counterterm. In the following calculations, we will use Eq. (40) as the physical and IR-finite energy flux operator.
@@ III.3 Local Hilbert space truncation at $j_{\rm max}=\frac{1}{2}$
The Hamiltonian density can be written as
@@
where the plaquette position is given by $(i,j)_{\rm center}=(1,1)$ and $(2,2)$ for $3\times 3$ and $5\times 5$ honeycomb lattices, respectively, as shown in Figs. 6 and 12.
@@ III.4 Quantum circuit for $j_{\rm max}=\frac{1}{2}$
We now discuss how to construct the quantum circuit for the $j_{\rm max}=\frac{1}{2}$ case. To prepare the ground state, we implement the adiabatic evolution starting from the electric ground state that corresponds to |$\ket{0}^{\otimes n}$ state [^8] In the qubit mapping $|0\rangle$ corresponds to the spin-down state, where $n=N_{i}N_{j}$ is the size of the honeycomb lattice as shown in Fig. 6. We apply the Trotter decomposition for adiabatic evolution, which approximates the full evolution operator as
$\mathcal{T}e^{-i\int_{0}^{t}{\rm d}t^{\prime}H_{\rm ad}(t^{\prime})}=\prod_{k=1}^{N_{T}}\prod_{(i,j)}e^{-i\Delta tH_{ij}^{\rm el}}e^{-i\frac{k}{N_{T}}\Delta tH_{ij}^{\rm mag}}e^{O(\Delta t)^{2}}\,,$ (47)
where $\mathcal{T}$ is the time-ordering operator and $H_{\rm ad}(t)=H^{\rm el}+\frac{t}{T}H^{\rm mag}$ with $T=N_{T}\Delta t$.
The implementation of the Hadamard test requires that the $\mathcal{O}$, $T_{0i}$ and real-time evolution operators are controlled by the ancilla qubit (see Fig. 4). A controlled-real-time-evolution operator $C[e^{-iH_{ij}t}]$ can be written as
@@
For the controlled-$T_{0,(i,j)\to(i^{\prime},j^{\prime})}$ operators, inspecting Eq. (III.3), we notice that the $T_{0,(i,j)\to(i^{\prime},j^{\prime})}$ operator is written as a sum of six operators ($\sigma^{z}_{ij}\sigma^{y}_{i^{\prime}j^{\prime}}D_{i^{\prime}j^{\prime}}$, $\sigma^{y}_{i^{\prime}j^{\prime}}D_{i^{\prime}j^{\prime}}$, $\sigma^{x}_{i^{\prime}j^{\prime}}\sigma^{y}_{ij}\widetilde{D}_{i^{\prime}j^{\prime}}D_{ij}$ and those with $i,j\leftrightarrow i^{\prime},j^{\prime}$), which can be implemented on quantum circuits in the diluted operator method. We note that the operators $D_{ij}$ and $\tilde{D}_{ij}$ lead to non-unitary evolution, while the rest can be decomposed into $\sigma^{x}$, $\sigma^{y}$ and $\sigma^{z}$ gates and thus can be implemented with standard unitary circuits. We compile $D_{ij}$ (and $\widetilde{D}_{ij}$) using the diluted operator method discussed in Sec. II.3, by adding an ancilla qubit and implementing the following operator
@@ IV.1.1 $3\times 3$ lattice
On a $3\times 3$ honeycomb lattice, as shown in Fig. 6, the local source operator given by the magnetic energy term in Eq. (46) is applied at the central plaquette at position $(1,1)$. We exactly diagonalize the Hamiltonian and calculate energy correlators according to Eqs. (II.2) and (12). Due to the small size of our lattice, we will only consider the first time integral step in Eq. (II.1.1). First, as a test, we study the integrand of the one-point energy correlator, i.e., the expectation value of the energy flux operator, as a function of time. Different $\vec{n}$s then pick out different directions of energy flows from the center $(1,1)$. We denote the expectation value of the energy flux operator from $(i,j)\to(i^{\prime},j^{\prime})$ at time $t$ as
@@
The results are shown in Fig. 8 for three directions: $(1,1)\to(1,2)$, $(1,1)\to(0,2)$ and $(1,1)\to(2,1)$. The $3\times 3$ honeycomb lattice shown in Fig. 6 has two reflection symmetries: with respect to the axis connecting $(0,0)$ and $(2,2)$ and that connecting $(0,2)$ and $(2,0)$. Since the local source operator is applied at $(1,1)$ and does not break either reflection symmetry, energy flows in the system still respect both reflection symmetries. As a result, the energy fluxes along $(1,1)\to(1,2)$ and $(1,1)\to(2,1)$ are exactly the same at all time. The system also respects rotational symmetry approximately till the perturbation created by the local source term reaches the boundary and bounces back, which is reflected in the approximate agreement between the energy flux along $(1,1)\to(0,2)$ and the other two energy fluxes in Fig. 8 for $t<13$. This rotational symmetry is related to the $N=1$ case of Eq. (2) and shows that the limit of the integral cannot exceed $t=13$ for one-point correlator.
Figure 8: Time dependence of energy flux operator on a $3\times 3$ honeycomb lattice with $ag^{2}=1$ and $j_{\rm max}=\frac{1}{2}$. The local source term is located at the central plaquette, i.e., $(1,1)$. The energy flux from $(1,1)$ to $(1,2)$ in black and that from $(1,1)$ to $(2,1)$ in red are exactly the same because of the reflection symmetry of the lattice along the axis connecting $(0,0)$ and $(2,2)$. The energy flux from $(1,1)$ to $(0,2)$ in blue agrees with the other two approximately up to $t=13$, which indicates the rotational symmetry is approximately respected at early time. At later time, perturbations reach the boundary and bounce back. The dynamics is influenced by the boundary effect, which breaks the rotational symmetry.
@@
We further test energy conservation in our lattice setup by computing the energy density at position $(1,1)$ and the integrated total energy flux going out of it. The energy density is calculated by taking the expectation value of the Hamiltonian density $H_{ij}$ defined in Eq. (III.3)
@@
where $i=1$, $j=1$ and the summation over $(i^{\prime},j^{\prime})$ includes $(0,1)$, $(0,2)$, $(1,0)$, $(1,2)$, $(2,0)$ and $(2,1)$. The results are depicted in Fig. 9. The energy density at $(1,1)$ keeps decreasing till $t\approx 12$, which is roughly the same time when the breaking of the rotational symmetry becomes manifest in Fig. 8. The energy density at $(1,1)$ starts to increase when the net energy flow bounces back from the boundary and flows inward towards the central plaquette. The energy decrease at position $(1,1)$ is exactly equal to the integrated total energy flux flowing out of it, as indicated by constant red dashed line in Fig. 9. This demonstrates energy conservation is maintained well in our lattice setup of the energy density and energy flux operators.
@@
Figure 11: Three-point energy correlator as a function of the coupling at $\theta_{12}=\theta_{23}=\theta_{31}=\frac{2\pi}{3}$ (black solid), compared with the magnitude of the one-point energy correlator cubed (blue dotted) on a $3\times 3$ honeycomb lattice with $j_{\rm max}=\frac{1}{2}$ and a local source at $(1,1)$. The vanishing of the one-point energy correlator around $ag^{2}=1$ is a lattice artifact, caused by the fact that the energy flux detectors are next to the local source. As a result, as the coupling decreases the net energy flux starts to flow inward and thus becomes negative in our convention.
Finally we study two-point and three-point energy correlators as functions of the coupling. As mentioned in Sec. II, it is important to maintain the spacelike separation between energy flux operators for the whole integration time domain by using the zigzag integration contour. However, on a $3\times 3$ lattice with the source placed at $(1,1)$, we do not have much space to move energy flux operators as time changes. So we just fix the position of the energy flux operator and only integrate for a short period of time, expecting that in such a short time period, the world lines of different energy flux operators will maintain spacelike separation. This requirement immediately excludes the configuration where two energy flux operators are on top of or next to each other.

### v2

@@ II.3.2 Hadamard test
After preparing the interacting ground state, we introduce an ancilla qubit and perform the Hadamard test: We first apply a Hadamard gate to an ancilla qubit, then apply a controlled-unitary gate for the operator insertions in the real-time correlator $\langle{\rm vac}|\mathcal{O}^{\dagger}T_{0i}(t_{1})T_{0j}(t_{2})\mathcal{O}|{\rm vac}\rangle$, [^5] We suppress the spatial dependence here for convenience. The time dependence is simulated as shown Eq. (II.2). and finally apply another Hadamard gate and measure the ancilla qubit, as shown in Fig. 4a The probability of the ancilla qubit being in $|0\rangle$ state is given by
@@
The source ($\mathcal{O}$), sink ($\mathcal{O}^{\dagger}$) and the energy flux $T_{0i}$ operators are not unitary, making their direct implementation in the quantum circuit hard. Here we assume the source and sink operators to be Hermitian (the energy flux operator $T_{0i}$ is Hermitian) and propose three different methods to implement them.
@@
    Taylor expansion of exponential. Suppose that $\mathcal{A}$ is a Hermitian operator and we have to obtain $\mathcal{A}\ket{\psi}$, we can implement two quantum circuits where one circuit implements $e^{i\theta\mathcal{A}}$ gate and the other $e^{-i\theta\mathcal{A}}$. The difference of the two circuits gives [^6] In practice, the difference is taken classically on the level of circuit measurement results rather than in the quantum circuit.
@@
    In this way, we can implement $\mathcal{A}\ket{\psi}$ up to a small error when $\theta$ is small. In our case, we have to implement four different non-unitary operators, $\{\mathcal{O}^{\dagger}=\mathcal{O},\,T_{0i}(t_{1}),\,T_{0j}(t_{2}),\,\mathcal{O}\}$. In order to obtain the real-time correlator $\langle{\rm vac}|\mathcal{O}T_{0i}(t_{1})T_{0j}(t_{2})\mathcal{O}|{\rm vac}\rangle$, we should run 16 different quantum circuits, in each of which we implement the operator of the form $e^{is_{1}\theta\mathcal{O}}e^{is_{2}\theta T_{0i}(t_{1})}e^{is_{3}\theta T_{0j}(t_{2})}e^{is_{4}\theta\mathcal{O}}$ for one sign combination $s_{i}=\pm$. There are a total of 16 sign combinations, corresponding to 16 quantum circuits. By taking a proper linear combination of the results from the 16 quantum circuits, we find
@@
    where the $2\times 2$ block structure is associated with the additional ancilla qubit. After the implementation of $U_{\mathcal{A}}$, the ancilla qubit is measured. When the measurement returns |$\ket{0}$, the correct operator is applied, i.e., $\mathcal{A}$. The success probability of applying $\mathcal{A}$ to |$\ket{\psi}$ is given by $P_{s}=\|\frac{\mathcal{A}}{\|\mathcal{A}\|}|\psi\rangle\|$, which is equal to the expectation value $\langle{\psi}|\frac{\mathcal{A}^{2}}{\|\mathcal{A}\|^{2}}|{\psi}\rangle$.
@@
    Linear combination of unitaries (LCU). The $\mathcal{O}$ and $T_{0i}$ operators of our interest can be written as sums of Pauli strings that are unitary operators. Therefore, we can apply the method of LCU [143]. To implement an operator $\mathcal{A}$ given by a linear combination of unitary operators $\mathcal{A}=\sum_{j=1}^{N}\kappa_{j}U_{j}$ with $\kappa_{j}$ real numbers, we can add $m$ additional ancilla qubits, where $m$ is given by $\lceil\log_{2}N\rceil$, and implement the quantum circuit shown in Fig. 5. The $V$ gate is given by a unitary that implements
@@
    Figure 5: LCU quantum circuit. The $V$ and $U_{\rm LCU}$ gates are defined in Eqs. (20) and (21), respectively.
    Analogous to the diluted operator method, the correct implementation of $\mathcal{A}$ is obtained if all the $m$ ancilla qubits are measured to be |$\ket{0}$. The success probability is given by
    $P_{s}=\frac{\|\sum_{j=1}^{N}\kappa_{j}U_{j}|\psi\rangle\|}{\|\kappa\|_{1}^{2}}\,.$ (22)
@@ III.1 Hamiltonian on honeycomb lattice
with $a,b,c\in\{1,2,3\}$ for SU(2) adjoint indices. $\varepsilon^{abc}$ represents the structure constant of the non-Abelian group and is the Levi-Civita tensor for the SU(2) case.
@@
Electric fields are generators of time-independent gauge transformation. On the honeycomb lattice, since electric fields live on links, they can generate gauge transformations on either red or black vertices. So we distinguish these two types by the subscript $R$ for red and $B$ for black. The gauge transformation is given by
@@
Because of the Peter-Weyl theorem [145], we can work in the electric basis where the basis states are specified by three quantum numbers $|jm_{R}m_{B}\rangle$, analogous to the total angular momentum quantum number $j$ and two third components $m_{R}$ and $m_{B}$ for the red and black vertices, respectively. In the electric basis, the matrix elements of the electric fields are given by [146, 147, 148]
@@
where $|j\rangle$ and $|J\rangle$ denote the states on a link where the electric field squared acts. The matrix elements of the electric energy are diagonal in the electric basis and they are independent of whether the red or black vertex is used. The matrix elements of the magnetic energy (the honeycomb plaquette) are given by [149, 150, 144] (see Refs. [151, 152, 153, 154, 155] for the square plaquette case)
@@ III.2 Energy flux operator
To construct the energy flow operator, we need specific components of the stress-energy tensor. In particular, the energy flux operator is given as
@@
which we need to integrate in order to build the energy flow operator as in Eq. (1). Here $k,\ell$ denote spatial indices. Discretized stress-energy tensor operators on square lattices have been constructed in Ref. [156].
On a honeycomb lattice, the position ${\boldsymbol{x}}$ of the $T_{0k}(\boldsymbol{x})$ operator is given by the red vertex as shown in Fig. 6, which can be labeled by $(i,j)$. The natural directions are no longer $\hat{x}$ and $\hat{y}$, but rather $\hat{e}_{1}$, $\hat{e}_{2}$ and $\hat{e}_{3}$. We can use the lattice discretized version of Eq. (36) and define three energy flux operators corresponding to the three electric field operators $E_{Ri}^{a}$ along the three directions $\hat{e}_{i}$ (see Appendix A). These energy fluxes point from red vertices to red vertices. Therefore, to indicate both the spatial position of the energy flux operator and the direction of the energy flow, we find it more convenient to introduce a notation to denote the energy flux from the red vertex $(i,j)$ pointing towards $(i^{\prime},j^{\prime})$ as
@@
From the expression given in Eq. (36), we can express $T_{0,(i,j)\to(i^{\prime},j^{\prime})}$ for our $2+1D$ SU(2) on the honeycomb lattice in terms of the electric field and the plaquette operator. We carry out this exercise in Appendix A. However, the energy flux operator obtained by discretizing Eq. (36) is a lattice bare operator. In general, one needs to tame UV divergences and renormalize them in the continuum limit, since the lattice breaks Lorentz symmetry and the stress-energy tensor obtained from discretizing the continuum expression may not satisfy the conservation equation $\partial_{\mu}T^{0\mu}=0$ on the lattice. As energy correlators describe how energy flows in the system, it is important to use a lattice energy flux operator that exactly conserves energy in the lattice calculation. Such a physical and IR-finite operator can be constructed by considering the Hamiltonian density at lattice position $(i,j)$ in the Heisenberg picture $H_{ij}(t)=e^{iHt}H_{ij}e^{-iHt}$. Taking time derivative gives
$\frac{{\rm d}H_{ij}(t)}{{\rm d}t}=i[H,H_{ij}(t)]\,.$ (38)
@@
On a honeycomb lattice, six possible energy flux directions $(i^{\prime},j^{\prime})$ for a given $(i,j)$ are $(i^{\prime},j^{\prime})\in\{(i,j+1),(i+1,j),(i+1,j-1),(i,j-1),(i-1,j),(i-1,j+1)\}$, as shown in Fig. 7, where the current $T_{0,(i,j)\to(i^{\prime},j^{\prime})}$ is nonvanishing. In Appendix A, we explicitly demonstrate that the lattice bare energy flux operator obtained by discretizing Eq. (36) is identical to Eq. (40) up to boundary terms and UV counterterm. In the following calculations, we will use Eq. (40) as the physical and IR-finite energy flux operator.
@@ III.3 Local Hilbert space truncation at $j_{\rm max}=\frac{1}{2}$
The Hamiltonian density per honeycomb plaquette can be written as
@@
where the plaquette position is given by $(i,j)_{\rm center}=(1,1)$ and $(2,2)$ on $3\times 3$ and $5\times 5$ honeycomb lattices, respectively, as shown in Figs. 6 and 12.
@@ III.4 Quantum circuit for $j_{\rm max}=\frac{1}{2}$
We now discuss how to construct the quantum circuit for the $j_{\rm max}=\frac{1}{2}$ case. To prepare the ground state, we implement the adiabatic evolution starting from the electric ground state that corresponds to |$\ket{1}^{\otimes n}$ state, [^9] In the qubit mapping $|0\rangle$ corresponds to the spin-up state while $|1\rangle$ represents the spin-down state. where $n=N_{i}N_{j}$ is the size of the honeycomb lattice as shown in Fig. 6. We apply the Trotter decomposition for adiabatic evolution, which approximates the full evolution operator as
$\mathcal{T}e^{-i\int_{0}^{t_{\rm ad}}{\rm d}t^{\prime}H_{\rm ad}(t^{\prime})}=\prod_{k=1}^{N_{\rm ad}}\prod_{(i,j)}e^{-i\Delta tH_{ij}^{\rm el}}e^{-i\frac{k}{N_{\rm ad}}\Delta tH_{ij}^{\rm mag}}e^{O(\Delta t)^{2}}\,,$ (47)
where $\mathcal{T}$ is the time-ordering operator and $H_{\rm ad}(t)=H^{\rm el}+\frac{t}{t_{\rm ad}}H^{\rm mag}$ with $t_{\rm ad}=N_{\rm ad}\Delta t_{\rm ad}$.
The implementation of the Hadamard test requires that the $\mathcal{O}$, $T_{0i}$ and real-time evolution operators are controlled by the ancilla qubit (see Fig. 4). A controlled-real-time-evolution operator $C[e^{-iHt}]$ can be written as
@@
For the controlled-$T_{0,(i,j)\to(i^{\prime},j^{\prime})}$ operators, inspecting Eq. (III.3), we notice that the $T_{0,(i,j)\to(i^{\prime},j^{\prime})}$ operator is written as a sum of six operators ($\sigma^{z}_{ij}\sigma^{y}_{i^{\prime}j^{\prime}}D_{i^{\prime}j^{\prime}}$, $\sigma^{y}_{i^{\prime}j^{\prime}}D_{i^{\prime}j^{\prime}}$, $\sigma^{x}_{i^{\prime}j^{\prime}}\sigma^{y}_{ij}\widetilde{D}_{i^{\prime}j^{\prime}}D_{ij}$ and those with $i,j\leftrightarrow i^{\prime},j^{\prime}$), which can be implemented in quantum circuits via the diluted operator method. The operators $D_{ij}$ and $\tilde{D}_{ij}$ lead to non-unitary evolution, while the rest can be decomposed into $\sigma^{x}$, $\sigma^{y}$ and $\sigma^{z}$ gates and thus can be implemented with standard unitary circuits. We compile $D_{ij}$ (and $\widetilde{D}_{ij}$) using the diluted operator method discussed in Sec. II.3, by adding an ancilla qubit and implementing the following operator
@@ IV.1.1 $3\times 3$ lattice
On a $3\times 3$ honeycomb lattice, as shown in Fig. 6, the local source operator given by the magnetic energy term in Eq. (46) is applied at the central plaquette at position $(1,1)$. We exactly diagonalize the Hamiltonian and calculate energy correlators according to Eqs. (II.2) and (12). First, as a test, we study the integrand of the one-point energy correlator, i.e., the expectation value of the energy flux operator, as a function of time. Different $\vec{n}$ vectors then pick out different directions of energy flows from the center $(1,1)$. We denote the expectation value of the energy flux operator from $(i,j)\to(i^{\prime},j^{\prime})$ at time $t$ as
@@
The results are shown in Fig. 8 for three directions: $(1,1)\to(1,2)$, $(1,1)\to(0,2)$ and $(1,1)\to(2,1)$. The $3\times 3$ honeycomb lattice shown in Fig. 6 has two reflection symmetries with respect to the axis connecting $(0,0)$ and $(2,2)$ and that connecting $(0,2)$ and $(2,0)$. Since the local source operator is applied at $(1,1)$ and does not break either reflection symmetry, energy flows in the system still respect both reflection symmetries. As a result, the energy fluxes along $(1,1)\to(1,2)$ and $(1,1)\to(2,1)$ are exactly the same at all time. The system also respects rotational symmetry approximately till the perturbation created by the local source term reaches the boundary and bounces back, which is reflected in the approximate agreement between the energy flux along $(1,1)\to(0,2)$ and the other two energy fluxes in Fig. 8 for $t<13$. This rotational symmetry is related to the $N=1$ case of Eq. (2) and shows that the limit of the integral cannot exceed $t=13$ for one-point correlator.
Figure 8: Time dependence of energy flux expectation values on a $3\times 3$ honeycomb lattice with $ag^{2}=1$ and $j_{\rm max}=\frac{1}{2}$. The local source term is located at the central plaquette, i.e., $(1,1)$. The energy flux from $(1,1)$ to $(1,2)$ in black and that from $(1,1)$ to $(2,1)$ in red are exactly the same because of the reflection symmetry of the lattice along the axis connecting $(0,0)$ and $(2,2)$. The energy flux from $(1,1)$ to $(0,2)$ in blue agrees with the other two approximately up to $t=13$, which indicates the rotational symmetry is approximately respected at early time. At later time, perturbations reach the boundary and bounce back. The dynamics is influenced by the boundary effect, which breaks the rotational symmetry.
@@
We further test energy conservation in our lattice setup by computing the energy density per plaquette at position $(1,1)$ and the integrated total energy flux going out of it. The energy density is calculated by taking the expectation value of the Hamiltonian density $H_{ij}$ defined in Eq. (III.3),
@@
where $i=1$, $j=1$ and the summation over $(i^{\prime},j^{\prime})$ includes $(0,1)$, $(0,2)$, $(1,0)$, $(1,2)$, $(2,0)$ and $(2,1)$. The results are depicted in Fig. 9. The energy density at $(1,1)$ keeps decreasing till $t\approx 12$, which is roughly the same time when the breaking of the rotational symmetry becomes manifest in Fig. 8. The energy density at $(1,1)$ starts to increase when the net energy flow bounces back from the boundary and flows inward towards the central plaquette. The energy decrease at position $(1,1)$ is exactly equal to the integrated total energy flux going out of it, as indicated by the constant red dashed line in Fig. 9. This demonstrates energy conservation is maintained well in our lattice setup of the energy density and energy flux operators.
@@
Figure 11: Three-point energy correlator as a function of the coupling at $\theta_{12}=\theta_{23}=\theta_{31}=\frac{2\pi}{3}$ (red solid), compared with the magnitude of the one-point energy correlator cubed (blue dotted) on a $3\times 3$ honeycomb lattice with $j_{\rm max}=\frac{1}{2}$ and a local source at $(1,1)$. The vanishing of the one-point energy correlator around $ag^{2}=1$ is a lattice artifact, caused by the fact that the energy flux detectors are next to the local source. As a result, as the coupling decreases the net energy flux starts to flow inward and thus becomes negative in our convention.
Finally we study two-point and three-point energy correlators as functions of the coupling. As mentioned in Sec. II, it is important to maintain the spacelike separation between energy flux operators for the whole integration time domain by using the zigzag integration contour. However, on a $3\times 3$ lattice with the source placed at $(1,1)$, we do not have much space to move energy flux operators as time changes. So we just consider the first time integral step in Eq. (II.1.1) by fixing the position of the energy flux operator and only integrating for a short period of time, expecting that in such a short time period, the world lines of different energy flux operators will maintain spacelike separation. This requirement immediately excludes the configuration where two energy flux operators are on top of or next to each other.